### v1

@@ 4 Applications
After establishing the theoretical foundations of PEMC and demonstrating its effectiveness in controlled environments, we now apply it to real-world financial scenarios where the complexity of models often renders standard variance reduction techniques difficult or unavailable. In this section, we examine PEMC in practice through two complex pricing problems. In particular, we examine variance swaps pricing under stochastic (local) volatility models (Dupire et al. 1994, Derman and Kani 1994, Gatheral 2006), and the pricing of swaptions under the Heath-Jarrow-Morton (HJM) framework (Carmona 2007). In both cases, the path-dependent nature of the contracts and the complexity of the model dynamics make traditional control variates difficult.
4.1 Variance Swaps Pricing
@@
Variance swaps are financial derivatives that enable investors to trade future realized volatility against current implied volatility (Bayer et al. 1999). Unlike traditional options, which derive their value from the underlying asset’s price, variance swaps are based on the variance of the asset’s returns over a specified period (Derman and Fantazzini 1999). This unique structure allows for pure exposure to volatility, making variance swaps valuable tools for risk management and speculative strategies (Gatheral 2006, Protter 2010).
In this subsection, we present an application of the PEMC framework in pricing a variance swaps, with the underlying asset paths assumed to be generated according to a Heston model (1) and a Stochastic Local Volatility (SLV) model (Ren et al. 2007), respectively. These models represent two distinct approaches to stochastic volatility modeling in practical derivatives pricing: the Heston model offers a parsimonious parametric form with relatively few parameters to calibrate, while the SLV model provides greater flexibility through a data-intensive, non-parametric specification of the volatility surface. By examining both cases, we demonstrate PEMC’s versatility across the spectrum of model complexity.
4.1.1 Variance swaps under a Heston model
Under the Heston model (1), and the procedure follows from the illustration in Section 2.5. Given $\bm{\theta}_{\text{model}}:=(r,\kappa,\eta,\delta,\rho)$, the asset price $\{S_{t}\}_{t}$ is simulated by Euler’s scheme with fixed $T$ and $\Delta t$:
$\nu_{t+\Delta t}$ $\leftarrow\kappa(\eta-\nu_{t})\Delta t+\delta\sqrt{\nu_{t}}\Delta W^{\nu}_{t},$
$S_{t+\Delta t}$ $\leftarrow S_{t}\exp\Big{(}\big{(}r-\tfrac{1}{2}\nu_{t+\Delta t}^{2}\big{)}\Delta t+\sqrt{\nu_{t+\Delta t}}\Delta W^{S}_{t}\Big{)}.$
where $(\Delta W^{S}_{t},\Delta W^{\nu}_{t})\stackrel{{\scriptstyle\text{i.i.d}}}{{\sim}}\sqrt{\Delta t}\cdot\mathcal{N}\Big{(}\mathbf{0},[\begin{smallmatrix}1&\rho\\ \rho&1\end{smallmatrix}]\Big{)}$. For each asset path $\{S_{t}\}_{t}$, the payoff is
$f_{\bm{\theta}_{\text{payoff}}}(\{S_{t}\}_{t}):=\sqrt{\frac{252\times\sum_{t=1}^{T/\Delta t}\log(S_{t}/S_{t-1})^{2}}{T/\Delta t}}-K.$ (12)
Note that due to the linear form of strike $K$ in the payoff, it suffices to effectively removes strike $K$ from training. We also conveniently omit any discount factor. The parameters space $\bm{\Theta}$, as well as the evaluation $\bm{\theta}$, is summarized in Table 4.
Table 3: Parameter Setup in Heston Model
mode $S_{0}$ $r$ $\nu_{0}$ $\kappa$ $\eta$ $\delta$ $\rho$ $T$, $\Delta t$
$N_{\text{train}}=3,000,000$ $[50,150]$ $[0.01,0.05]$ $[0.1,0.375]^{2}$ $[1.5,4.5]$ $[0.1,0.3]^{2}$ $[0.1,1.0]$ $[-0.2,-0.9]$ $1$, $1/252$
evaluation $100$ $0.02$ $0.25^{2}$ $3.0$ $0.2^{2}$ $0.6$ $-0.4$
The feature $\bm{X}$ is constructed by
which are jointly distributed as $\sqrt{T}\cdot\mathcal{N}(0,[\begin{smallmatrix}1&\rho\\ \rho&1\end{smallmatrix}])$ just as in Section 2.5. For the prediction model training in PEMC algorithm 1, a basic Residual Network (ResNet)-style architecture (He et al. 2016) is adopted and trained on $N_{\text{train}}=3\times 10^{6}$ samples. The network consists of two Multi-Layer Perceptron (MLP) blocks connected through skip connections. The first MLP block processes $(\bm{\theta},\bm{X})$ together and pass them through hidden layers with dimension 512, eventually producing a 256-dimensional intermediate representation. This output is then fed into the second MLP block, which maintains the same hidden dimension of 256 throughout its layers. The complete architecture specification is provided in Table 4. At evaluation stage (i.e., Algorithm 2), we use $\bm{\theta}$ in Table 4 and set $N=10n$ as before.
Table 4: Hyper-parameter Setup for Neural Network
1st MLP block 2nd MLP Block
hidden dim: 512; output dim: 256 hidden dim: 256
Following our evaluation methodology from Section 3.5, we first establish a ground truth value by averaging $5\times 10^{7}$ MC samples. We then implement both standard MC and PEMC estimators (with $N=10n$) using sample sizes $n\in\{1000,2000,4000,8000,10000,20000\}$, repeating each experiment 200 times to assess sampling variability. As shown in Table 5, PEMC achieves a 30-50% reduction in root mean squared error compared to standard MC, consistent with the variance reduction levels we observed in Section 3.5. This demonstrates that PEMC’s effectiveness as a variance reduction technique carries over to more complex models and derivative products.
Table 5: RMSE from 200 Experiments for Pricing Variance Swap under the Heston Model
Method $n=1000$ $n=2000$ $n=4000$ $n=6000$ $n=8000$ $n=10000$ $n=20000$
Monte Carlo (MC) 0.0996 0.0659 0.0517 0.0391 0.0366 0.0290 0.0218
PEMC 0.0599 0.0387 0.0263 0.0233 0.0213 0.0179 0.0126
4.1.2 Variance swaps under an SLV model
Having demonstrated PEMC’s effectiveness for the Heston model in Section 4.1.1, we turn to the more sophisticated Stochastic Local Volatility (SLV) framework. While parametric models like Heston offer mathematical elegance and computational efficiency, they often lack the flexibility to fully capture market dynamics (Gatheral 2006). SLV models emerged as a hybrid approach that combines the market-implied local volatility surface with stochastic volatility dynamics (Guyon and Henry-Labordère 2014), providing practitioners with greater calibration flexibility and more accurate price reproduction across strike-maturity ranges (Ren et al. 2007).
@@ 4.1.2 Variance swaps under an SLV model
The increased modeling power of SLV, however, comes with significant computational challenges. It is known that SLV has no closed-form formula for even vanilla options (Tataru and Fisher 2010). Moreover, unlike the Heston model’s parsimonious parameter set, SLV requires handling a full volatility surface discretized on a dense 2D grid, effectively turning our parameter space $\bm{\Theta}$ for PEMC into a high-dimensional object. To efficiently process this grid-structured volatility data, we adopt a Convolutional Neural Network (CNN) architecture (LeCun et al. 1998) for our PEMC implementation, effectively feeding the local volatility surface as an image into PEMC. The CNNs are particularly well-suited for this task as they naturally exploit the spatial relationships in the volatility surface, similar to their success in image processing tasks (He et al. 2016, Krizhevsky et al. 2012).
@@
with $\langle\mathrm{d}W_{t}^{S},\mathrm{d}W_{t}^{\nu}\rangle=\rho$ and $\eta_{t}:=-\frac{\delta^{2}}{2\kappa}(1+e^{-2\kappa t}).$ Here $\sigma(\cdot,\cdot):\mathbb{R}\times\mathbb{R}^{+}\to\mathbb{R}^{+}$ is 2D function representing the local volatility surface and $\exp(\nu_{t})$ is a stochastic multiplier with $\exp(\nu_{0})=1$ and $\mathbb{E}[e^{2\nu_{t}}]=1$. In practice, local volatility surfaces are calibrated to and stored as discrete two-dimensional grids indexed by asset prices (spot) and time. During simulation, these discrete values are interpolated as needed to obtain volatilities at arbitrary price-time points (Coleman et al. 2001). For our PEMC implementation, we treat this discrete grid as part of our input parameter $\bm{\theta}_{\text{model}}$, reflecting how the market-calibrated surfaces would be used in practice. While local volatility models (Dupire et al. 1994, Derman and Kani 1994) and their calibration to market data constitute an extensive research area in their own right, our focus here is on the PEMC implementation where, a calibrated local volatility has been given for evaluation, regardless of the calibration method used to obtain it.
Following practical conventions but also simplifying for the sake of illustration, we store $\sigma(\cdot,\cdot)$ on a $|\mathcal{S}|\times|\mathcal{T}|$ grid, where $\mathcal{S}$ contains $|\mathcal{S}|$ equally-spaced price points in $[S_{\text{surface}}^{\min},S_{\text{surface}}^{\max}]$ and $\mathcal{T}$ contains $|\mathcal{T}|$ equally-spaced time points in $[t_{\text{surface}}^{\min},t_{\text{surface}}^{\max}]$. At each point on the grid, we store the value of local volatility according to Carmona (2007):
@@
for $p_{0}=0.3,p_{1}=0.5,p_{2}=0.2,\tau_{0}=0.4,\tau_{1}=0.3,\tau_{2}=0.6$ as in Figure 2 in Carmona (2007). While this analytical form (15) is used in our data-generating process, it is important to note that PEMC treats the surface as any market-calibrated volatility surface - accessing it only through its discrete grid values. When sampling $\bm{\theta}\sim\bm{\Theta}$ to produce $\sigma$ grid, we add a $\mathcal{N}(0,\xi^{2})$ noise independently to all the $|\mathcal{S}|\times|\mathcal{T}|$ points in the grid, on top of their baseline value $\sigma_{base}$ (15). A path is then generated by Euler’s scheme:
@@
with $(\Delta W^{S}_{t},\Delta W^{\nu}_{t})\stackrel{{\scriptstyle\text{i.i.d}}}{{\sim}}\sqrt{\Delta t}\cdot\mathcal{N}(\mathbf{0},[\begin{smallmatrix}1&\rho\\ \rho&1\end{smallmatrix}])$, and $\widehat{\sigma}(S_{t},t)$ obtained by interpolations of the grid $\sigma$. The input $(\bm{\theta},\bm{X})$ can be represented as
$\text{feature}_{i}=\{\underbrace{\{\sigma^{(s,t)}\}_{s\in\mathcal{S},t\in\mathcal{T}}}_{\text{surface info }},\underbrace{r,\delta,\kappa,\rho,\mu}_{\bm{\theta}_{\text{model}}},\underbrace{S_{0},\nu_{0}}_{\bm{\theta}_{\text{simulation}}},\underbrace{K}_{\bm{\theta}_{\text{payoff}}},\underbrace{(W^{S}_{T},W^{\nu}_{T})}_{\mathbf{X}(\bm{\theta})}\}.$
The parameters space $\bm{\Theta}$, as well as the evaluation $\bm{\theta}$, is summarized in Table 6.
Table 6: Parameter Setup in SLV Model
@@
To handle the 2D grid of high-dimensional volatility surface data, we design a two-branch neural network architecture, which is illustrated in Figure 1. The first branch processes the discretized volatility surface $\sigma^{2}(x,t)$ using a CNN architecture inspired by VGG (Simonyan and Zisserman 2014), which has become a standard choice for image processing tasks and is readily available in modern deep learning packages like PyTorch (Paszke et al. 2019). This branch then consists of two 2D convolutional layers interspersed with ReLU activations, followed by a MaxPool2d operation for dimensionality reduction. The surface features are then flattened through a fully connected layer to produce an embedding. The second branch handles the remaining model parameters in $(\bm{\theta},\bm{X})$ through a series of fully connected layers with dropout regularization, batch normalization, and ReLU activations, ultimately producing another embedding. Finally, the two separate embeddings are then fed into a “Synthesizer” module, which combines the information through additional fully connected layers with dropout and ReLU activations to produce the final prediction. This architecture choice is motivated by the proven effectiveness of CNNs in handling grid-structured data (LeCun et al. 1998, He et al. 2016), and particularly the VGG architecture’s success in extracting hierarchical features while maintaining relative simplicity (Simonyan and Zisserman 2014). The details of NN architecture are summarized in Table 7.
Figure 1: Neural Network architecture for the SLV Model.
Table 7: Neural Network Architecture Parameters
@@
Using the same experimental setup as before, we evaluate PEMC’s performance in the SLV setting across sample sizes $n\in\{1000,2000,4000,8000,10000,20000\}$ with $N=10n$, benchmarking against a ground truth computed from $5\times 10^{7}$ MC samples. The results, based on 200 independent experiments and shown in Table 8, reveal that PEMC’s effectiveness persists even in this more complex setting. Despite the added complexity of handling high-dimensional volatility surfaces, PEMC achieves a 30-40% reduction in mean squared error compared to standard MC, demonstrating its robustness as a variance reduction technique across different model frameworks.
Table 8: RMSE from 200 Experiments for Pricing Variance Swap under the SLV Model
@@
4.2 Swaptions in HJM Models
Interest rate derivatives play a crucial role in financial markets, with swaptions being particularly important instruments for managing interest rate risk (Brigo and Mercurio 2006). A swaption gives its holder the right to enter into an interest rate swap at a future date, providing flexibility in hedging future interest rate exposures (Hull 2018). The pricing of these instruments, however, presents significant computational challenges due to the high-dimensional nature of interest rate modeling (Andersen and Piterbarg 2010). In this subsection, we demonstrate PEMC’s application to swaption pricing under the Heath-Jarrow-Morton (HJM) framework (Heath et al. 1992). The HJM model directly describes the evolution of the entire forward rate curve, offering greater flexibility than traditional short-rate models. For illustration purposes, we focus on a one-factor specification with exponential volatility structure (Glasserman 2013b), though the framework readily extends to multi-factor cases. This model choice represents a balance between analytical tractability and practical relevance, allowing us to demonstrate PEMC’s effectiveness in a high-dimensional setting while maintaining computational feasibility.
4.2.1 Swaptions and HJM Model
A swaption is a contract granting its holder the right, but not the obligation, to enter into an interest rate swap at a future date. In a standard interest rate swap, one party agrees to pay a fixed rate while receiving a floating rate, and the other party does the opposite. Consider a swap with $n_{p}$ fixed payment periods, each of length $\Delta t^{\prime}$, starting at time $t^{\prime}_{0}$ and ending at time $t^{\prime}_{n_{p}}=t^{\prime}_{0}+\sum_{l=1}^{n_{p}}\Delta t^{\prime}$. The value of this swap at time $t^{\prime}_{0}$ is:
@@ 4.2.1 Swaptions and HJM Model
where $C$ is the notional amount (contract size), $R$ is the fixed rate, and $B(t^{\prime}_{0},t^{\prime}_{l})$ is the discount factor from $t^{\prime}_{0}$ to $t^{\prime}_{l}$. A swaption provides the holder with the option to enter into this swap at $t^{\prime}_{0}$. The payoff of the swaption is simly
@@
where $\mu(t,T)$ is the drift, $\sigma(t,T)$ is the volatility function of the forward rate, and $W(t)$ is a Brownian motion. In contrast to short-rate models (e.g., Vasicek (Vasicek 1977) or Cox-Ingersoll-Ross (Cox et al. 1985)), which only model the dynamics of the short-term interest rate, the HJM model directly models the dynamics of the entire term structure of interest rates (Heath et al. 1992). The HJM model is widely used in practice because of its flexibility in modeling interest rate derivatives like swaptions and its ability to incorporate complex volatility structures (Brigo and Mercurio 2006, Andersen and Piterbarg 2010). However, the model’s generality also leads to the need for sophisticated numerical methods for simulation Glasserman (2013b). A key property of the HJM model is the no-arbitrage condition (Glasserman 2013b), which specifies the drift completely by the volatility:
@@
4.2.2 PEMC for HJM
Just as with the local volatility surface case, in practice $\sigma(t,T)$ cannot be predefined parametrically and must be calibrated from market data of caps, floors, and swaptions, yielding a discrete grid of values. However, for demonstration purposes, we employ a classical exponential decay specification as our baseline model inspired from Glasserman (2013b):

### v2

@@ 5. Applications
After establishing the theoretical foundations of PEMC and demonstrating its effectiveness in controlled environments, we now apply it to real-world financial scenarios where the complexity of models often renders standard variance reduction techniques difficult or unavailable. In this section, we first examine PEMC in practice through two complex options pricing problems. In particular, we examine variance swaps pricing under stochastic local volatility models [46, 43, 56], and the pricing of swaptions under the Heath-Jarrow-Morton (HJM) framework [34]. In both cases, the path-dependent nature of the contracts and the complexity of the model dynamics make traditional control variates difficult.
5.1. Variance Swaps in stochastic local vol models
@@
Variance swaps are financial derivatives that enable investors to trade future realized volatility against current implied volatility [14]. Unlike traditional options, which derive their value from the underlying asset’s price, variance swaps are based on the variance of the asset’s returns over a specified period [42]. This unique structure allows for pure exposure to volatility, making variance swaps valuable tools for risk management and speculative strategies [56, 108]. In this subsection, we focus on using PEMC to price variance swaps under a stochastic local volatility (SLV) model [111], as closed-form formulas exist for variance swaps in stochastic volatility models such as Heston [134], but not for SLV [125] which provides greater flexibility through a data-intensive, non-parametric specification of the volatility surface [56]. SLV models emerged as a hybrid approach that combines the market-implied local volatility surface with stochastic volatility dynamics [63], providing practitioners with greater calibration flexibility and more accurate price reproduction across strike-maturity ranges [111].
@@
Unlike the Heston model’s parsimonious parameter set, SLV requires handling a full volatility surface discretized on a dense 2D grid, effectively making the parameter space $\boldsymbol{\Theta}$ for PEMC high-dimensional. To efficiently process this grid-structured volatility data, we adopt a Convolutional Neural Network (CNN) architecture [83] for PEMC, effectively feeding the local vol surface as an image into predictive modeling. The CNNs are particularly well-suited for this task as they naturally exploit the spatial relationships in the volatility surface, similar to their success in image processing tasks [64, 81].
@@
with $\langle\mathrm{d}W_{t}^{S},\mathrm{d}W_{t}^{\nu}\rangle=\rho dt$ and $\eta_{t}:=-\frac{\delta^{2}}{2\kappa}(1+e^{-2\kappa t}).$ Here $\sigma(\cdot,\cdot):\mathbb{R}\times\mathbb{R}^{+}\to\mathbb{R}^{+}$ is 2D function representing the (interpolated) local volatility surface and $\exp(\nu_{t})$ is a stochastic multiplier with $\exp(\nu_{0})=1$ and $\mathbb{E}[e^{2\nu_{t}}]=1$. In practice, local volatility surfaces are calibrated to and stored as discrete two-dimensional grids indexed by asset prices (spot) and time. During simulation, these discrete values are interpolated as needed to obtain volatilities at arbitrary price-time points [37]. For our PEMC implementation, we treat this discrete grid as part of our input parameter $\boldsymbol{\theta}_{\text{model}}$, reflecting how the market-calibrated surfaces would be used in practice. While local volatility models [46, 43] and their calibration to market data constitute an extensive research area in their own right, our focus here is on the PEMC implementation. Thus, we assume a calibrated local volatility has been given for evaluation, regardless of the method used to obtain it. Following practical conventions, we store $\sigma(\cdot,\cdot)$ on a $|\mathcal{S}|\times|\mathcal{T}|$ grid, where $\mathcal{S}$ contains $|\mathcal{S}|$ equally-spaced price points in $[S_{\text{surface}}^{\min},S_{\text{surface}}^{\max}]$ and $\mathcal{T}$ contains $|\mathcal{T}|$ equally-spaced time points in $[t_{\text{surface}}^{\min},t_{\text{surface}}^{\max}]$. At each point on the grid, we store the value of local volatility according to [34]:
@@
for $p_{0}=0.3,p_{1}=0.5,p_{2}=0.2,\tau_{0}=0.4,\tau_{1}=0.3,\tau_{2}=0.6$ as in Figure 2 in [34]. While this analytical form (14) is used in our data-generating process, it is important to note that PEMC treats the surface as any market-calibrated volatility surface - accessing it only through its discrete grid values. When sampling $\boldsymbol{\theta}\sim\boldsymbol{\Theta}$ to produce $\sigma$ grid, we add a $\mathcal{N}(0,\xi^{2})$ noise independently to all the $|\mathcal{S}|\times|\mathcal{T}|$ points in the grid, on top of their baseline value $\sigma_{base}$ (14). A path is then generated by Euler’s scheme:
@@
with $(\Delta W^{S}_{t},\Delta W^{\nu}_{t})\stackrel{{\scriptstyle\text{i.i.d}}}{{\sim}}\sqrt{\Delta t}\cdot\mathcal{N}(\mathbf{0},[\begin{smallmatrix}1&\rho\\ \rho&1\end{smallmatrix}])$, and $\widehat{\sigma}(S_{t},t)$ obtained by interpolations of the grid $\sigma$. The input $(\boldsymbol{\theta},\boldsymbol{X})$ can be represented as
$\text{feature}_{i}=\{\underbrace{\{\sigma^{2}(s,t)\}_{s\in\mathcal{S},t\in\mathcal{T}}}_{\text{surface info }},\underbrace{r,\delta,\kappa,\rho,\mu}_{\boldsymbol{\theta}_{\text{model}}},\underbrace{S_{0},\nu_{0}}_{\boldsymbol{\theta}_{\text{simulation}}},\underbrace{K}_{\boldsymbol{\theta}_{\text{payoff}}},\underbrace{(W^{S}_{T},W^{\nu}_{T})}_{\mathbf{X}(\boldsymbol{\theta})}\}.$
The parameters space $\boldsymbol{\Theta}$, as well as the evaluation $\boldsymbol{\theta}$, is summarized in Table 1.
Table 1. Parameter Setup in SLV Model
@@
To handle the 2D grid of high-dimensional volatility surface data, we design a two-branch neural network architecture, which is illustrated in Figure 2. The first branch processes the discretized volatility surface $\sigma^{2}(x,t)$ using a CNN architecture inspired by VGG [118], which has become a standard choice for image processing tasks and is readily available in modern deep learning packages like PyTorch [103]. This branch then consists of two 2D convolutional layers interspersed with ReLU activations, followed by a MaxPool2d operation for dimensionality reduction. The surface features are then flattened through a fully connected layer to produce an embedding. The second branch handles the remaining model parameters in $(\boldsymbol{\theta},\boldsymbol{X})$ through a series of fully connected layers with dropout regularization, batch normalization, and ReLU activations, ultimately producing another embedding. Finally, the two separate embeddings are then fed into a “Synthesizer" module, which combines the information through additional fully connected layers with dropout and ReLU activations to produce the final prediction. This architecture choice is motivated by the proven effectiveness of CNNs in handling grid-structured data [83, 64], and particularly the VGG architecture’s success in extracting hierarchical features while maintaining relative simplicity [118]. The details of NN architecture are summarized in Table 2.
Figure 2. Neural Network architecture for the SLV Model.
Table 2. Neural Network Architecture Parameters
@@
Finally, we evaluate PEMC’s performance in the SLV setting across sample sizes $n\in\{1000,2000,$ 4000,8000,10000,20000}4000,8000,10000,20000\}4000 , 8000 , 10000 , 20000 } with $N=10n$, benchmarking against a ground truth computed from $5\times 10^{7}$ MC samples. The results, based on 200 independent experiments and shown in Table 3, reveal that PEMC’s effectiveness persists even in this more complex setting. Despite the added complexity of handling high-dimensional volatility surfaces, PEMC achieves a 30-40% reduction in mean squared error compared to standard MC, demonstrating its robustness as a variance reduction technique across different model frameworks.
Table 3. RMSE from 200 Experiments for Pricing Variance Swap under the SLV Model
@@
5.2. Swaptions in HJM Models
Interest rate derivatives play a crucial role in financial markets, with swaptions being particularly important instruments for managing interest rate risk [29]. A swaption gives its holder the right to enter into an interest rate swap at a future date, providing flexibility in hedging future interest rate exposures [73]. The pricing of these instruments, however, presents significant computational challenges due to the high-dimensional nature of interest rate modeling [5]. In this subsection, we demonstrate PEMC’s application to swaption pricing under the Heath-Jarrow-Morton (HJM) framework [65]. The HJM model directly describes the evolution of the entire forward rate curve, offering greater flexibility than traditional short-rate models. For illustration purposes, we focus on a one-factor specification with exponential volatility structure [60], though the framework readily extends to multi-factor cases. A swaption is a contract granting its holder the right, but not the obligation, to enter into an interest rate swap at a future date. In a standard interest rate swap, one party agrees to pay a fixed rate while receiving a floating rate, and the other party does the opposite. Consider a swap with $n_{p}$ fixed payment periods, each of length $\Delta t^{\prime}$, starting at time $t^{\prime}_{0}$ and ending at time $t^{\prime}_{n_{p}}=t^{\prime}_{0}+\sum_{l=1}^{n_{p}}\Delta t^{\prime}$. The value of this swap at time $t^{\prime}_{0}$ is:
@@ 5.2. Swaptions in HJM Models
where $C$ is the notional amount (contract size), $R$ is the fixed rate, and $B(t^{\prime}_{0},t^{\prime}_{l})$ is the discount factor from $t^{\prime}_{0}$ to $t^{\prime}_{l}F$. A swaption provides the holder with the option to enter into this swap at $t^{\prime}_{0}$. The payoff of the swaption is simply
@@
where $\mu(t,T)$ is the drift, $\sigma(t,T)$ is the volatility function of the forward rate, and $W(t)$ is a Brownian motion. In contrast to short-rate models (e.g., Vasicek [127] or Cox-Ingersoll-Ross [39]), which only model the dynamics of the short-term interest rate, the HJM model directly models the dynamics of the entire term structure of interest rates [65]. The HJM model is widely used in practice because of its flexibility in modeling interest rate derivatives like swaptions and its ability to incorporate complex volatility structures [29, 5]. However, the model’s generality also leads to the need for sophisticated numerical methods for simulation [60]. A key property of the HJM model is the no-arbitrage condition [60], which specifies the drift completely by the volatility:
@@
5.2.1. PEMC for HJM
Just as with the local volatility surface case, in practice $\sigma(t,T)$ cannot be predefined parametrically and must be calibrated from market data of caps, floors, and swaptions, yielding a discrete grid of values. However, for demonstration purposes, we employ a classical exponential decay specification as our baseline model inspired from [60]: